OpenStudy (anonymous):

Show that the function f given by \(f(\mathbf{x}) = |\mathbf{x}|\) is continuous on \(\mathbb{R}^n\)

OpenStudy (amistre64):

hmmm thats vector notation isnt it

OpenStudy (anonymous):

Yes, x would be an n-dimensional vector

OpenStudy (amistre64):

then |x| = sqrt(x1^2+x2^2+...+xn^2)

OpenStudy (anonymous):

I had that, from there would it be ok to just take the limit to a, and substitute it in since the square root is continuous?

OpenStudy (amistre64):

if the right and left limits, or in this case the disk, ball, something, limits exist and the value of f(x) is equal to the limit at x, then it is continuous right?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

might want to determine if there is any special handling at the "origin" since it might be hard to express that as a onesided limit

OpenStudy (amistre64):

we could go: d = sqrt(....); and d^2 = (.....) maybe

OpenStudy (zarkon):

\[\bigg|\|x\|-\|y\|\bigg| \leq \|x-y\|\]

OpenStudy (zarkon):

I would do an \(\epsilon,\delta\) proof using the reverse triangle inequality (above)

OpenStudy (anonymous):

Alright I'll give that a shot :) If I run into trouble I'll tell you guys :D